Solve for x: 5x-4(x-1)=-3(2x-5)-x
Answer by checkley71(8403)

5X-4(X-1)=-3(2X-5)-X
5X-4X+4=-6X+15-X
5X-4X+6X+X=15-4
8X=11
X=11/8
PROOF
5*11/8-4(11/8-1)=-3(2*11/8-5)-11/8
55/8-44/8+4=-3(22/8-5)-11/8
11/8+4=-3(22/8-40/8)-11/8
11/8+32/8=-3(-18/8)-11/8
43/8=54/8-11/8
43/8=43/8